supposing there are 1.50 x 10^6 g of sulfur dioxide emitted by an industrial plant , how many moles of calcium carbonate and oxygen that are needed to remove the given amount of pollutant ? spacecrafts and submarines use LiOH to capture exhaled CO2 in the reaction producing lithium carbonate and water vapor .​
1
Expert's answer
2020-12-03T14:06:54-0500

2SO2 + 2CaCO3 + O2 → 2CaSO4 + 2CO2

​M(SO2) = 64.06 g/mol

n(SO2) "= \frac{m}{M} = \frac{1.50 \times 10^6}{64.06} = 23.41 \times 10^3 \; mol"

n(CaCO3) "= n(SO_2) = 23.41 \times 10^3 \; mol"

n(O2) "= \frac{1}{2}n(SO_2) = \frac{23.41 \times 10^3}{2} = 11.70 \times 10^3 \; mol"

Answer: "23.41 \times 10^3 \; mol \;and \;11.70 \times 10^3 \; mol"